According to the U.S. Census Bureau, among males over the age of 24, 13.4% did not complete high school, 31.9% completed high school, and 16.5% attended some college without graduating. (Source: U.S. Department of Commerce, Census Bureau, Current Population Survey Data on Educational Attainment.) a. Compute and interpret the probability that a randomly selected male in the United States who is over the age of 24 did not attend college.
Answer #1

Probability = P(Did not complete high school )+ P( Completed high school)

= 0.134 + 0.319 = 0.453

Means 45.3% of males over 24 did not attend the college